OpenStudy (anonymous):

HELP!! For certain category hurricanes, the cube of the diameter D (in miles) of the hurricane is roughly proportional to the square of the hurricane's duration t (in hours). Write an equation to represent this relationship, then solve your equation for D.

OpenStudy (unklerhaukus):

\[D^3\quad\widetilde\propto\quad t^2\]

OpenStudy (nubeer):

D^3 ∝ t^2 D^3 = kt^2 D = (kt^2)^1/3

OpenStudy (anonymous):

Thank you!